Circular No. 6745

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUNGRAZING COMETS C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery of three more probable Kreutz sungrazing comets. Measurements by St. Cyr and D. Biesecker have been reduced by G. V. Williams and are given on MPEC 1997-S05, 1997-S06, and 1997-S07, together with tentative orbital elements by B. G. Marsden. C/1997 R1, discovered by Biesecker, appeared in C3 coronagraphic data only and brightened to at best m_{1} ∼ 8. C/1997 R2, also discovered by Biesecker, appeared in both C3 and C2 data and brightened to m_{1} ∼ 6.5. C/1997 R3, discovered by S. Stezelberger, was also in both C3 and C2 data and peaked at m_{1} ∼ 6. None of the comets showed any evidence of a tail.

1996 UT  α_{2000}  δ_{2000}  Comet
Sept. 8.149  10^{h}47.^{m}4  + 4^{o}43′  C/1997 R1
14.841  11 09.1  + 2 04  C/1997 R2
15.998  11 10.1  + 1 36  C/1997 R3

SUPERNOVA 1997df IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (R = 19) on V and R CCD images taken on Aug. 15 by S. Chan at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997df is located near Abell 3223 at α = 4^{h}09^{m}13^{s}.8, δ = −30^{o}42′23′′.3 (equinox 2000.0), which is 5′′.5 west and 2′′.9 north of the host galaxy. A nearby star has position end figures 16^{s}.17, 41′17′′.6. SN 1997df was not visible on similar images taken on Aug. 3 (limiting mag R = 21), but is present in V and R images taken on Aug. 27, Sept. 8, Sept. 10, and Sept. 17.

COMET C/1997 J2 (MEUNIER-DUPOUY) Total visual magnitude estimates: Aug. 9.24 UT, 12.2 (C. E. Spratt,

Victoria, BC, 0.20-m reflector); 25.81, 11.4 (V. S. Nevski, Vitebsk, Belarus, 0.30-m reflector); 30.91, 11.7 (J. Carvajal, Soria, Spain, 0.32-m reflector); Sept. 5.81, 12.0 (K. Sárneczky, Ráktanya, Hungary, 0.30-m reflector); 11.09, 11.8 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 16.79, 11.7 (M. Plsek, Lelekovice, Czech Republic, 0.20-m reflector); 22.04, 12.0 (J. Bortle, Stormville, NY, 0.41-m reflector).

1997 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green